# PRISM-Loc: a Lightweight Long-range LiDAR Localization in Urban Environments with Topological Maps

### v1

@@ 5 $P_{u}\leftarrow upsample(P_{t},f=3)$;
[…]
@@ V EXPERIMENTS
V-A Setup and Dataset
We evaluated the proposed method and the competitors on the outdoor part of the ITLP-Campus dataset [29]. This dataset contains several runs of a mobile differential-drive Husky robot on a university campus. The length of each run varied from 3 to 4 km and from 1 to 1.5 hours. These runs were recorded in different seasons and daytimes, and all of them cover the whole campus.
For all the methods, we used track 03 for map creation, and tracks 00, 01, and 02 for evaluation of the localization within the built map. The most challenging are the tracks 00 and 01 which were recorded in winter (whereas track 03 was recorded in spring). In addition to the dynamic changes between the tracks (parked and moving cars, pedestrians, etc.), the tracks 00 and 01 contain snow piles with heights up to 1 m.
@@ V-B Baselines
The first one is a classical point cloud based approach, where the algorithm operates with a point cloud itself. For localization in such methods, an accurate and dense point cloud should be given as an input. GLIM [30] was used for creating a 3D map. Unfortunately, we cannot use GLIM localization, because it is a proprietary solution. Instead of it, we used HDL_Localization method [31] to localize within the GLIM-built map.
GLIM method uses fixed‐lag smoothing combined with a keyframe‐based point cloud matching approach to robustly estimate sensor trajectories. In addition, multi‐camera visual constraints are incorporated to further enhance stability and accuracy. The authors used the GPU-accelerated scan matching process, which enables the system to directly minimize registration errors between submaps in a densely connected factor graph – circumventing some of the limitations of conventional Gaussian pose graph optimization – while still achieving real-time performance despite its high computational load.
Another method was BEVPlace++ [25]. It introduces a fast, robust, and lightweight LiDAR global localization method designed for unmanned ground vehicles. The approach begins by projecting sparse LiDAR point clouds into bird’s-eye view (BEV) images, which serve as a compact and informative representation. A specially designed Rotation Equivariant and Invariant Network (REIN) then extracts rotation-equivariant local features and rotation-invariant global descriptors from these BEV images. This enables a two-stage localization process where coarse place recognition is performed via global descriptors and refined through local feature matching with RANSAC for accurate 3-DoF pose estimation. The method’s efficiency, combined with its robustness to view and sensor variations without requiring dense pose supervision, makes it an appealing baseline for real-time global localization tasks.
@@
To overcome such limitations, two filtering strategies were devised, both leveraging the ”last known database position index” – a concept predicated on the assumption that sequential database indices correspond to roughly proximate poses. This approach circumvents reliance on external localization modalities (e.g., odometry) by exploiting temporal continuity and the known approximate inter-frame distance between consecutive database entries.
The first filter (denoted as ”BEVPlace++” in results) enforces temporal coherence by constraining the search space to database indices incrementally adjacent to the last confirmed position. Candidate poses retrieved via FAISS similarity search are evaluated based on their index proximity to the prior validated index. While this method improves robustness against isolated outliers, it does not account for loop closure events, leading to potential track loss.
TABLE I: Results on ITLP-Campus dataset
Method Mean ATE, m Median ATE, m $SR_{loc}$ RAM usage, GB Map size, GB Localization time, s
GLIM + HDL loc. 3.6* 3.5* 0.3 11 + 1.2** 0.4 3.0
BEVPlace++ 14.4 11.0 0.84 15.5 15.3 1.0
BEVPlace++ (outlier filtering) 20.3 6.8 0.95 15.5 15.3 1.0
PRISM-Loc 8.2 0.7 0.98 4.1 0.2 0.5
* ATE was estimated only before the localization failure
** 11 GB of RAM was used by GLIM for map creation, and 1.2 GB of RAM was used by HDL for localization
Filter 2 (denoted as ”BEVPlace++ wo outliers) incorporates spatial distance thresholds between the current pose estimate and the last validated pose. To mitigate cumulative drift caused by consecutive outliers, a buffer of recent outlier indices and their associated positional offsets is maintained. When a given distance threshold is exceeded, the system records the indices and distances of the outliers. If number of consecutive outliers have been scored, the filter pulls out the index with the minimum distance and makes it the last index. However, this method requires careful calibration of two hyperparameters: the size of the emission buffer and the threshold distance, which directly affect the robustness to stable emission sequences.
Two baseline frameworks were evaluated against the proposed method: (1) HDL localization based on GLIM created map, and (2) BEVPlace++ with 2 outlier filtering algorithms, which were developed for rough real-time pose estimation.
@@ V-C Results
The results are shown in Table I. The comparison of the estimated trajectories with all the methods and the boxplots for trajectory errors are shown in Fig. 7. At each track, HDL method failed to localize in the GLIM’s map after traveling about 1 km. The sparse voxelized point cloud map for the 3 km route weighed only about 400 MB; however, each localization procedure in it took about 3 s, with complete localization loss in the middle of the track. Because of significant changes between the tracks, the pure BEVPlace++ method without filtering exhibited about 16% of place recognition errors, which led to a mean ATE of 20 m (see Fig. 6). The suggested filtering approach eliminated most of the place recognition errors, however, the percent of localization fails remained high. The map for BEVPlace++ took 15 GB on disk storage and in the RAM because of storing large local descriptors for each location. Also, the localization pipeline took in average 1.0 sec (about 700 ms for place recognition, and 300 ms for pose estimation using a local descriptor).
The proposed PRISM-Loc method made localization losses and errors in only 2% of cases thanks to our combined approach to localization. The median ATE reached the value of 0.7 m, significantly outperforming all the competitors. With such localization quality, our approach remained fast and lightweight – the map of all the campus consumed only 200 MB (the largest part of 4 GB RAM consumption is the reserving allocation for PyTorch models, this memory is not utilized, so memory usage could be optimized in the future). Also, the global localization time took only 0.5 s (0.1 s for place recognition, and 0.08 s for scan matching with each of five found candidates).
(a) Trajectory on $XY$ plane
(b) Box Plot
Figure 7: Comparison analysis for track 00
@@ VI CONCLUSION AND FUTURE WORK
We propose PRISM-Loc – a lightweight and robust method for localization in large outdoor environments using topological maps. The proposed method combines tracking through rough odometry estimations, a novel robust scan-matching technique integrated with the original curb detection algorithm, and place recognition for global location search. Experiments conducted on a challenging multi-season dataset covering a 3 km route demonstrate that the proposed method achieves robust localization, with a median trajectory error of 0.7 meters, despite significant object-wise and seasonal differences between the run and the map. Furthermore, PRISM-Loc maintains low memory usage and computational time throughout the entire route. Moving forward, we plan to enhance our approach by incorporating local pose filtering to improve robustness and to conduct extensive research on scan-matching techniques.

### v2

@@ 6 $P_{u}\leftarrow upsample(P_{t},f=3)$;
7 $mask_{1}\leftarrow P_{u}.z<\text{curb\_height}$ ;
8 $mask_{2}\leftarrow abs(P_{u}.z)>3\cdot\text{curb\_height}$;
@@ V EXPERIMENTS
V-A Setup and Datasets
We evaluated the proposed method and the competitors on three large-scale outdoor datasets. The first of them is the outdoor part of the ITLP-Campus dataset [melekhin2024itlp]. This dataset contains several runs of a mobile differential-drive Clearpath Husky robot on a university campus. These runs were recorded in different seasons and daytimes, and all of them cover the whole campus. The sensor data include scans from a 16-ray Velodyne LiDAR, images from two cameras, and wheel encoders’ data.
The second dataset includes the first 6 sequences of Oxford RobotCar Radar [barnes2020oxford], which were collected from an autonomous car in an urban environment. It contains data from two 32-ray LiDARs, IMU, and two cameras. The summary of used datasets is shown in Table II.
TABLE II: Datasets used in the experiments
Dataset N of seqs Environment
Avg. track
length, km
IMU
ITLP-Campus [melekhin2024itlp] 4 Campus 3.3 No
RobotCar [barnes2020oxford] 6 Urban 7.8 Yes
For both datasets, we used one track for map creation, and the rest of the tracks were used for evaluation of the localization within the built map. For ITLP-Campus dataset, we created the map by track 03, and ran localization on track 00, 01, and 02. The main challenge of evaluation on this dataset is the seasonal difference between tracks (tracks 00 and 01 were recorded in winter and have a large amount of snow piles on lidar scans, and tracks 02 and 03 were recorded in spring). For Oxford RobotCar, we created the map by track 2019-01-11-14-02-26, and evaluated the localization methods on five other tracks recorded in January 2019. The main challenge of this dataset is the presence of a large amount of dynamic objects (like cars and pedestrians), which may obstruct lidar scan matching.
@@ V-B Baselines
The first one is a classical point cloud-based approach, where the algorithm operates with a point cloud itself. For localization in such methods, an accurate and dense point cloud should be given as an input. To create a 3D point cloud map, DLO [chen2022direct] method was used. It consists of two components, odometry estimation and 3D mapping. It cannot work in localization-only mode (only in SLAM mode), so, for localization in the built map, we used HDL_Localization method [koide_Portable_2019].
DLO is a lightweight 3D LiDAR-inertial SLAM method. For odometry estimation, it uses NanoGICP (an optimized version of Generalized ICP [koide2021voxelized] algorithm). Global 3D map is built as a set of 3D submaps, each submap is created from a single voxelized LiDAR scan. Such solution allows to create compact 3D maps for large environments. To build precise map relying on ground truth positions with DLO, we used these positions instead of DLO odometry estimations.
Another method was BEVPlace++ [luo_BEVPlace_2024]. It introduces a fast, robust, and lightweight LiDAR global localization method designed for unmanned ground vehicles. The approach begins by projecting sparse LiDAR point clouds into bird’s-eye view (BEV) images, which serve as a compact and informative representation. A specially designed Rotation Equivariant and Invariant Network (REIN) then extracts rotation-equivariant local features and rotation-invariant global descriptors from these BEV images. This enables a two-stage localization process where coarse place recognition is performed via global descriptors and refined through local feature matching with RANSAC for accurate 3-DoF pose estimation. The method’s efficiency, combined with its robustness to view and sensor variations without requiring dense pose supervision, makes it an appealing baseline for real-time global localization tasks.
@@
As a possible improvement, we consider a lightweight temporal-coherence filtering strategy that leverages the “last known database position index”—assuming sequential database indices correspond to nearby poses. This option reduces reliance on external localization modalities (e.g., odometry) by exploiting temporal continuity and the approximate inter-frame distance between consecutive database entries.
TABLE III: Localization quality values
Method ITLP-Campus RobotCar
Mean ATE Median ATE $SR_{loc}$ Mean ATE Median ATE $SR_{loc}$
GLIM + HDL loc. failed 0.04 4.3* 1.9* 0.72
BEVPlace++ 14.4 11.0 0.84 14.9 1.6 0.94
[ours] PRISM-Loc 0.5 0.3 0.99 5.9 1.8 0.90
* Localization failed on tracks 2019-01-15-13-06-37 and 2019-01-18-15-20-12;
ATE reported only up to failure
TABLE IV: Localization performance values
Method ITLP-Campus RobotCar
RAM, GB Map size, GB Runtime, s RAM, GB Map size, GB Runtime, s
DLO + HDL loc. 1.2 0.05 1.0 4.2 0.12 1.0
BEVPlace++ 1.5 0.8 0.4 2.6 10.6 0.4
[ours] PRISM-Loc 0.2 0.02 0.3 0.7 0.02 0.3
Concretely, the filter constrains the search space to indices adjacent to the last confirmed position: FAISS-retrieved candidates are re-scored by their proximity (in index space) to the prior validated index. This increases robustness to isolated outliers, though it does not handle loop closures and may still lead to track loss.
Two baseline frameworks were evaluated against the proposed method: (1) HDL localization based on a DLO-created map, and (2) BEVPlace++ (optionally augmented with the temporal-coherence filter) developed for rough real-time pose estimation.
@@ V-C Results
The results are shown in Table III. The comparison of the estimated trajectories with all the methods and the boxplots for trajectory errors are shown in Fig. 7. At each track of ITLP-Campus, HDL method failed to localize in the 3D map after traveling about 200 m. The sparse voxelized point cloud map for the 3 km route weighed only about 50 MB; however, each localization procedure in it took about 1 s, with complete localization loss in the beginning of the track. On dense urban environment of Oxford RobotCar dataset, HDL failed on two tracks of five, reaching 72% success rate, and 4 m mean ATE on succeeded tracks.
Because of significant changes between the tracks, the pure BEVPlace++ method without filtering exhibited about 16% of place recognition errors, which led to a mean ATE of 20 m (see Fig. 6). The suggested filtering approach eliminated most of the place recognition errors, however, the percent of localization fails remained high. The map for BEVPlace++ took 15 GB on disk storage and in the RAM because of storing large local descriptors for each location. Also, the localization pipeline took in average 1.0 sec (about 700 ms for place recognition, and 300 ms for pose estimation using a local descriptor).
Figure 7: Trajectory on XYXY plane
The proposed PRISM-Loc method made localization losses and errors in only 1% of cases on ITLP and in 10% of cases on Oxford RobotCar thanks to our combined approach to localization. The mean ATE on ITLP-Campus the value of 0.5 m, significantly outperforming all the competitors. With such localization quality, our approach remained fast and lightweight – the maps for both ITLP and Oxford RobotCar consumed only 20 MB (the largest part of 4 GB RAM consumption is the reserving allocation for PyTorch models, this memory is not utilized, so memory usage could be optimized in the future). Also, the global localization time took only 0.3 s (0.1 s for place recognition, and 0.04 s for scan matching with each of five found candidates).
@@ VI CONCLUSION AND FUTURE WORK
We propose PRISM-Loc – a lightweight and robust method for localization in large outdoor environments using topological maps. The proposed method combines tracking through rough odometry estimations, a novel robust scan-matching technique integrated with the original curb detection algorithm, and place recognition for global location search. Experiments conducted on two challenging large-scale datasets with route lengths up to 8 km demonstrate that the proposed method achieves robust localization, with a mean trajectory error of 0.5 meters, despite significant object-wise and seasonal differences between the run and the map. Furthermore, PRISM-Loc maintains low memory usage and computational time throughout the entire route. Moving forward, we plan to enhance our approach by incorporating local pose filtering to improve robustness and to conduct extensive research on scan-matching techniques.